OpenStudy (anonymous):

log_{5}(2x+3)=log_{5}3 answer is 16/5 how do i get to that answer? whats the process?

jimthompson5910 (jim_thompson5910):

Since the logs have the same base, you can say 2x+3 = 5

OpenStudy (anonymous):

because log is a one to one function, so \[\log(x)=\log(y)\iff x=y\]

jimthompson5910 (jim_thompson5910):

oops meant say 2x+3 = 3

OpenStudy (anonymous):

which means there is a mistake here somewhere right?

OpenStudy (anonymous):

since \(2x+3=3\implies x=0\)